Evaluate log 0.2 base of 25.
Answer by stanbon(75887)

Evaluate log 0.2 base of 25.
------
log25(0.2) = log(0.2)/log(25) = -0.5
----
Why??
log(0.2) = log(5^-1)
Log(25) = log(5^2)
----
Therefore:: log25(0.2) = -log(5)/2log(5) = -1/2 = -0.5
---------------
Cheere,
Stan H.
--------------